# What Does Neuro Mean to Cardio? Investigating the Role of Clinical Specialty Data in Medical LLMs

### v3

@@ Abstract
In this paper, we introduce S-MedQA, an English medical question-answering (QA) dataset for benchmarking large language models (LLMs) in fine-grained clinical specialties. S-MedQA has over 2020k examples, covers $15$ medical specialties, and QA pairs can have multiple specialty annotations (e.g., when a question is cross-disciplinary), constructed with both machine and expert verification to maximize data availability. We use S-MedQA to investigate the role of clinical specialty data in the knowledge-intensive scenario of medical QA. Our results show that 1) training on data from a clinical specialty does not necessarily lead to best performance on that specialty, and 2) regardless of the specialty the LLM was fine-tuned on, token probabilities of clinically relevant terms increase consistently across all specialties. Thus, we hypothesize improvement gains are derived mostly from domain shifting (e.g., general to medical) rather than specialty-specific knowledge injection, and suggest rethinking the role of fine-tuning data in the medical domain. To motivate further advancements in the clinical NLP field, we release S-MedQA and all code needed to reproduce all our experiments to the research community. [^1] https://anonymous.4open.science/r/S-MedQA-85FD
@@ 1 Introduction
To address this gap, we propose S-MedQA, the first English medical QA dataset with multiple clinical specialty annotations (see Figure 1 for an overview, and §2 for details). We build S-MedQA based on the widely used MedQA (Jin et al., 2021) and MedMCQA (Pal et al., 2022) datasets, incorporating both machine and expert verification to semi-automatically map samples onto clinical specialties at scale, while guaranteeing the quality of the annotations. We first use multiple prompts and a majority voting mechanism to label QA pairs with a single specialty. S-MedQA includes 15 specialties, each with hundreds to thousands of samples, and single specialty expert validation shows up to 97.8% accuracy in this annotation (see §2 for details). In order to account for cross-disciplinary questions, the second step in our data collection is to expand single-specialty annotations into multiple specialties. To do that, we use a multi-label conformal prediction procedure with coverage guarantees. We annotate a held-out set of QA pairs with medical experts and build a conformal multi-label annotation procedure that leads to a precision of $0.69$, recall of $0.52$, 2424% exactly correct predictions, and an average number of labels of $1.55$.
Figure 1: Overview of S-MedQA’s construction process. For single specialty annotation of each sample, we generate predictions using 5 different prompts and only keep those where predictions agree ($3+$, $4+$, or 55 times). A medical expert manually annotates S-MedQA’s entire validation and test set. We randomly sample $1,000$ questions from our train set and ask multiple medical experts to evaluate GPT-3.5’s predictions (we detail inter-annotator agreement in §2.2), achieving accuracies ranging from 90.890.897.897.8%. For multi-specialty annotation, we leverage conformal prediction to assign labels across multiple clinical specialties, using a held-out set of $300$ samples manually annotated by medical experts as calibration/test set. It achieves a precision of $0.69$, recall of $0.52$, with $24\%$ exactly correct matches and an average prediction length of $1.55$.
@@
Finally, we ask how do clinically-relevant terms for one specialty change before and after fine-tuning the LLM? We address this question in §3.4. We again use the clinical terms curated in §3.3 and analyze changes in term probabilities before and after fine-tuning on data from the same vs. different clinical specialties. Our results suggest that performance gains are driven more by domain shifts (from general to medical domain) than by fine-grained clinical knowledge injection.
@@
  • We introduce S-MedQA, a medical QA dataset spanning 15 clinical specialties with 4040k QA pairs annotated with high quality and human validated specialty annotations.
@@ 2.3 From single to multiple clinical specialties
We curate a held-out set of 300 samples, annotated independently by a medical expert (300 samples) and three medical students (100 samples each). We use 200 examples (66.6%) as a calibration set, and report results on the remaining 100 examples (33.3%).
@@ 3.1 Experimental setup
In all subsequent analyses below, we use the 6 most common specialties (Cardiology, Gastroenterology, Infectious diseases, Neurology, Obstetrics and Gynecology, and Pediatrics), with test set sample size of $80$, $83$, $102$, $74$, $88$, and $90$, respectively.
@@ 3.2 Cross-specialty evaluation
We experiment with 8 open-source LLMs: Llama2-Chat-7B, Llama2-Chat-13B, Llama-3.1-8B-Instruct, Llama-3.2-3B-Instruct (Touvron et al., 2023), Mistral-Instruct-v0.1, Mistral-Instruct-v0.2 (Jiang et al., 2023) , Bio-Medical-LLaMA-3-8B (Con, 2024), and OLMo-2-1124-7B-Instruct (OLMo et al., 2024). These include six general-purpose models (Llama-2, Llama-3, and Mistral-based models), one biomedical Llama model, and a fully open-source model (OLMo).
@@ Results.
Table 4 shows the performance of Llama-3.1-8B-Instruct fine-tuned independently on each specialty and tested on all six specialty test sets, as well as after fine-tuning on a combination of all six specialties. Are improvements truly indicative of knowledge acquisition or injection? Notably, for models fine-tuned on individual specialties, none of the best performing models were trained on the corresponding specialty’s training data, e.g., the best performance on the Cardio test set (69.9%) was achieved by the model trained on Neuro. If the improvements were due to knowledge injection, we would expect to see best-performing models consistently along the diagonal (e.g., Cardio $\rightarrow$ Cardio, Infect $\rightarrow$ Infect, and so on). We report results for other LLMs in §A.5 and note that similar inconsistencies hold, whereas with different transfer patterns across specialties. This inconsistency suggests that the performance gains from fine-tuning primarily arise from factors other than knowledge injection, partially supporting Zhou et al. (2024)’s hypothesis in the clinical domain.
@@ 4 Discussion
In knowledge-intensive domains such as clinical specialties, the role of fine-tuning remains underexplored. Clinical questions across specialties often pose substantial knowledge barriers, and the capabilities of current LLMs on specialty-specific tasks are still unclear. Moreover, it remains uncertain whether these tasks benefit more from explicit knowledge injection or from effective knowledge transfer within the base model. This paper benchmarks these real-world questions in a fine-grained manner and provides preliminary evidence to explain the role of fine-tuning in the clinical domain.
@@ 6 Conclusions
Experiments on knowledge transfer patterns
@@
Language and healthcare system
@@ A.5 Additional cross-specialty evaluation results
In Tables 56,  7,  4,  9,  10,  11 we show cross-specialty evaluation matrices for Llama2-7b-chat, Llama2-13b-chat, Mistral-7b-instruct-v0.1, Llama-3.1-8B-Instruct, Llama-3.2-3B-Instruct, Bio-Medical-LLaMA-3-8B, and OLMo-2-1124-7B-Instruct in addition to our main results (in §3.2). Here we also observe that the best performance on each per-specialty test set is not achieved by the model that is tuned on training data from the same specialty (i.e., we observe this in $\sim$70% of the cases).
@@
A.6 High-level Concepts in SNOMED-CT
The high-level SNOMED-CT disorder concepts we use when mapping questions to clinical specialities are: cardiology (Disorder of cardiovascular system (SCTID: 49601007)), gastroenterology (Disorder of digestive system (SCTID: 53619000)), infectious diseases (Infectious disease (SCTID: 40733004)), obstetrics (Disorder of female reproductive system (SCTID: 363124003), Disorder of fetus and/or mother during labor (SCTID: 1269083008)), neurology (Disorder of nervous system(SCTID: 118940003)), pediatrics (Behavioral and emotional disorder with onset in childhood (SCTID: 231538003), Disorder of fetus and/or newborn (SCTID: 414025005), Developmental disorder (SCTID: 5294002)).
A.7 Token probability change

### v4

@@ Abstract
In this paper, we introduce S-MedQA, an English medical question-answering (QA) dataset for benchmarking large language models (LLMs) in fine-grained clinical specialties. S-MedQA has over 2424k examples, covers $15$ medical specialties, and QA pairs can have multiple specialty annotations, e.g., when a question is cross-disciplinary, constructed with both machine and expert verification to maximize data availability and reliability. We use S-MedQA to investigate the role of clinical specialties in the knowledge-intensive scenario of medical QA. Our results show that 1) training on data from a clinical specialty does not necessarily lead to best performance on that specialty, and 2) regardless of the specialty the LLM was fine-tuned on, token probabilities of clinically relevant terms increase consistently across all specialties. Thus, we hypothesize improvement gains, at least in our settings, are derived mostly from domain shifting (e.g., general to medical) rather than specialty-specific knowledge injection, and suggest rethinking the role of fine-tuning data in the medical domain. To motivate further advancements in the clinical NLP field, we release S-MedQA and all code needed to reproduce all our experiments to the research community. [^1] https://github.com/nlp4health-lab/S-MedQA
@@ 1 Introduction
To address this gap, we propose S-MedQA, the first English medical QA dataset with multiple clinical specialty annotations (see Figure 1 for an overview, and §2 for details). We build S-MedQA based on the widely used MedQA (Jin et al., 2021) and MedMCQA (Pal et al., 2022) datasets, incorporating both machine and expert verification to semi-automatically map samples onto clinical specialties at scale, while guaranteeing the quality of the annotations. We first use multiple prompts and a majority voting mechanism to label QA pairs with a single specialty. S-MedQA includes 15 specialties, each with hundreds to thousands of samples, and single specialty expert validation shows up to 97.8% accuracy in this annotation (see §2 for details). In order to account for cross-disciplinary questions, the second step is to expand single-specialty annotations into multiple specialties. We use a multi-label conformal prediction procedure with coverage guarantees. We annotate a held-out set of QA pairs with medical experts with a Jaccard Similarity of $0.82$, and build a conformal multi-label annotation procedure that leads to a precision of $0.69$, recall of $0.52$, 2424% exactly correct predictions, and an average number of labels of $1.55$.
Figure 1: Overview of S-MedQA’s construction process. For single specialty annotation of each sample, we generate predictions using 5 different prompts and only keep those where predictions agree ($3+$, $4+$, or 55 times). A medical expert manually annotates S-MedQA’s entire validation and test set. We randomly sample $1,000$ questions from our train set and ask multiple medical experts to evaluate GPT-3.5’s predictions (we detail inter-annotator agreement in §2.2), achieving accuracies ranging from 90.890.897.897.8%. For multi-specialty annotation, we leverage conformal prediction to assign labels across multiple clinical specialties, using a held-out set of $300$ samples manually annotated by the same medical expert and three medical students as calibration/test set, with a Jaccard Similarity of 0.82, reflecting a high inter-annotation agreement. It achieves a set precision of $0.69$, set recall of $0.52$, with $24\%$ exactly correct matches and an average prediction length of $1.55$.
@@
Finally, we ask how does the probability of clinically-relevant terms for one specialty change before and after fine-tuning the LLM? We address this question in §3.4. We again use the clinical terms curated in §3.3 and analyze changes in term probabilities before and after fine-tuning on data from the same vs. different clinical specialties. Our results suggest that performance gains are driven more by domain shifts (from general to medical) than by fine-grained clinical knowledge injection.
@@
  • We introduce S-MedQA, a medical QA dataset spanning 15 clinical specialties with 2424k QA pairs annotated with high quality and human validated specialty annotations.
@@ 2.3 From single to multiple clinical specialties
We curate a held-out set of 300 samples, annotated independently by a medical expert (300 samples) and three medical students (100 samples each). We measure the Jaccard Similarity between the expert and the students for the inter-annotator agreement, with a similarity score of 0.82. We use 200 examples (66.7%) as a calibration set, and report results on the remaining 100 examples (33.3%).
@@ 3.1 Experimental setup
In all subsequent analyses below, we use the 6 most common specialties (Cardiology, Gastroenterology, Infectious diseases, Neurology, Obstetrics and Gynecology, and Pediatrics). The number of test samples in each specialty is $80$, $83$, $102$, $74$, $88$, and $90$, respectively. The final number of samples in the test set is $400$, $415$, $510$, $370$, $440$, and $450$ (i.e., times 5 shuffles).
@@ 3.2 Cross-specialty evaluation
We experiment with 8 open-source LLMs: Llama2-Chat-7B, Llama2-Chat-13B, Llama-3.1-8B-Instruct, Llama-3.2-3B-Instruct (Touvron et al., 2023), Mistral-Instruct-v0.1, Mistral-Instruct-v0.2 (Jiang et al., 2023) , Bio-Medical-LLaMA-3-8B (ContactDoctor, 2024), and OLMo-2-1124-7B-Instruct (OLMo et al., 2025). These include six general-purpose models (Llama-2, Llama-3, and Mistral-based models), one biomedical Llama model, and a fully open-source model (OLMo).
@@ Results.
Table 4 shows the performance of Llama-3.1-8B-Instruct fine-tuned independently on each specialty and tested on all six specialty test sets, as well as after fine-tuning on a combination of all six specialties. Are improvements truly indicative of knowledge acquisition or injection? Notably, for models fine-tuned on individual specialties, none of the best performing models were trained on the corresponding specialty’s training data, e.g., the best performance on the Cardio test set (69.9%) was achieved by the model trained on Neuro. If the improvements were due to knowledge injection, we would expect to see best-performing models consistently along the diagonal (e.g., Cardio $\rightarrow$ Cardio, Infect $\rightarrow$ Infect, and so on). We report results for other LLMs in §A.5 and note that similar inconsistencies hold, whereas with different transfer patterns across specialties. This pattern suggests that performance gains from fine-tuning primarily arise from factors other than knowledge injection, partially supporting Zhou et al. (2024)’s hypothesis in the clinical domain.
The number of per-specialty examples used to train Llama-3.1-8B-Instruct is imbalanced (e.g., see Table 1). To tackle this imbalance, we randomly select 913 samples from each of the six specialties’ training data, which corresponds to the least number of samples among the six specialties. We fine-tune Llama-3.1-8B-Instruct on this data and we show the results in §A.6. We also conduct 20 bootstrap resampling experiments on the test sets using the Llama-3.1-8B-Instruct model fine-tuned on the whole (i.e., imbalanced) training data to tackle the issue of limited testing data. We randomly sample the same number of samples as the test sizes from the original test sets with replacement. We show the results for this experiment in §A.6. Both experiments show that most of the best performing models on each per-specialty’s test set is not the same model tuned on the training data from the same specialty. This demonstrates the robustness of our main results, i.e., best models are almost always off-diagonal.
@@ 4 Discussion
In knowledge-intensive domains such as healthcare, the role of fine-tuning remains underexplored. Clinical questions across specialties often pose substantial knowledge barriers, and the capabilities of current LLMs on specialty-specific tasks are still unclear. Moreover, it remains uncertain whether these tasks benefit more from explicit knowledge injection or from effective knowledge transfer within the base model. This paper benchmarks these real-world questions in a fine-grained manner and provides preliminary evidence to explain the role of fine-tuning in the clinical domain.
@@ 6 Conclusions
Experiments on knowledge transfer patterns.
@@
Language and healthcare system.
@@ A.5 Additional cross-specialty evaluation results
In Tables 4567910, and 11 we show cross-specialty evaluation matrices for Llama2-7b-chat, Llama2-13b-chat, Mistral-7b-instruct-v0.1, Llama-3.1-8B-Instruct, Llama-3.2-3B-Instruct, Bio-Medical-LLaMA-3-8B, and OLMo-2-1124-7B-Instruct in addition to our main results (in §3.2). Here we also observe that in $\sim$70% of the cases the best performance on each per-specialty test set is not achieved by the model that is tuned on training data from the same specialty.
@@
A.6 Robustness of cross-specialty evaluation
We carry out 2 experiments to show the robustness of our cross-specialty evaluation results in Table 4. We show the results of our cross specialty evaluation fine-tuned on balanced training set sizes in Table 12, and the results of the 20 bootstrap resampling experiments on the test sets in Table 13. Both experiments show that most of the best performing models on each per-specialty’s test set is not the same model tuned on the training data from the same specialty. This demonstrates the robustness of our main results, i.e., best models are almost always off-diagonal.
Test Sets Cardio Gastro Infect Neuro Obstetrics Pediatrics avg.
Train Sets Llama-3.1 64.0 67.5 65.3 66.8 65.0 64.5 65.5
Cardio 71.8 77.4 72.3 76.1 69.9 71.0 73.0
Gastro 69.6 77.1 72.4 77.2 69.1 71.9 72.8
Infect 68.8 74.5 73.3 77.4 69.8 71.3 72.4
Neuro 70.8 77.1 73.4 76.9 68.6 72.1 73.0
Obstetrics 69.8 75.0 72.9 75.4 68.7 69.4 71.9
Pediatrics 68.9 75.3 72.8 77.0 69.1 70.4 72.2
Table 12: Cross specialty evaluation results fine-tuned on training data with equal sizes.
Test Sets Cardio Gastro Infect Neuro Obstetrics Pediatrics avg.
Train Sets Llama-3.1 70.1 72.3 71.0 72.8 70.2 69.4 71.0
Cardio 79.7 75.4 72.3 77.6 73.1 71.1 73.6
Gastro 70.0 77.4 72.2 77.2 73.0 72.7 73.6
Infect 70.4 73.8 72.7 77.3 71.0 70.3 72.6
Neuro 72.0 76.0 72.7 77.3 72.0 71.3 73.4
Obstetrics 70.3 75.4 71.9 75.7 72.5 69.5 72.4
Pediatrics 71.2 74.8 72.9 76.4 71.4 70.8 72.8
Combined 63.1 81.4 77.5 78.2 73.3 76.6 76.7
Table 13: Cross specialty evaluation results of 20 resamplings of the test sets.
A.7 High-level Concepts in SNOMED-CT
The high-level SNOMED-CT disorder concepts we use when mapping questions to clinical specialities are: cardiology (Disorder of cardiovascular system, SCTID: 49601007), gastroenterology (Disorder of digestive system, SCTID: 53619000), infectious diseases (Infectious disease, SCTID: 40733004), obstetrics (Disorder of female reproductive system, SCTID: 363124003; Disorder of fetus and/or mother during labor, SCTID: 1269083008), neurology (Disorder of nervous system, SCTID: 118940003), pediatrics (Behavioral and emotional disorder with onset in childhood, SCTID: 231538003; Disorder of fetus and/or newborn, SCTID: 414025005; Developmental disorder, SCTID: 5294002).
A.8 Token probability change